OpenStudy (anonymous):

Given the function g(x) = −2x + 4, find g(−5). −6 4.5 6 14

OpenStudy (anonymous):

need help n explain it to me thxs

OpenStudy (anonymous):

Simply put -5 in for x and solve.

OpenStudy (anonymous):

plug-in x = -5 g(x) = −2x + 4, find g(−5). g(-5) = -2(-5) + 4 g(-5) = 10 + 4 g(-5) = 14